A group of students washed cars to raise money. The net amount A, in dollars, raised by washing k cars is given by the function A(K) = 4K-30 .If the group washed 15 cars, what is the net amount they raised?
(A) $10
(B) $15
(C) $20
(D) $25
(E) $30

Answer by palanisamy(496)

Given, A(K) = 4K-30
Plug K = 15
We get A(15) = 4*15-30=60-30 = $ 30